OpenStudy (anonymous):

Find the area of the region between the loops of the limacon r=5/2 + 5cos(theta).

OpenStudy (zarkon):

use \[\frac {1}{2}\int\limits_a^b \left[r(\theta)\right]^2\, d\theta\] twice